OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 3x

OpenStudy (anonymous):

Is this the correct answer? 2 cos^2x sin x + sin x - 2 sin^3x

OpenStudy (anonymous):

use sin(a+b) write sin(2x+x) then use formula

OpenStudy (anonymous):

which gives you sin2xcosx+cos2xsinx :)

OpenStudy (anonymous):

finally sin3x=3sinx-4(sinx)^3

OpenStudy (anonymous):

now you can wtire sin2x = sin(x+x) and cos2x= cos(x+x) apply the formula for them you will get all terms in form of sin x and cos x :D

hartnn (hartnn):

2 cos^2x sin x + sin x - 2 sin^3x is also correct. but that can be simplified using cos^2 x = 1-sin^2 x and then your answer will only be in sin x

OpenStudy (anonymous):

Thanks for your help guys!

OpenStudy (anonymous):

^^ welcome